# One Shot Learning as Instruction Data Prospector for Large Language Models

### v3

@@ 4.1 Experimental Setup
We exploit the Alpaca dataset [41]. It is an important resource in the open-source community for instruction tuning, which is constructed by employing the self-instruct [45] method to distill instruction data from text-davinci-003. The success of this dataset in fine-tuning the LLaMA model has sparked a series of explorations into instruction fine-tuning [23, 20, 54, 28, 64, 12].
@@
The predefined task set under consideration is associated with the earlier cited task pool, which is used to compute golden scores for instructions. These data primarily serve to assess the model’s capacity to generalize across a variety of tasks, allowing for considerable leeway in data selection. The predefined task set is deemed adequate provided it includes a sufficient quantity of data and encapsulates a wide array of tasks. Given that the Alpaca dataset naturally displays these characteristics, we have randomly selected 1,000 examples from it to form the predefined task set.
@@
Numerous endeavors are dedicated to evaluating the abilities of LLMs [27, 14, 13, 17, 26, 62]. However, it remains an ongoing and unresolved issue. Presently, a commonly adopted strategy involves employing the robust language model GPT-4 as the evaluator [26, 62, 8, 7]. In this framework, we employ two methods to assess the model’s capabilities. The first involves rating the responses generated by the model on a scale of 1 to 10. A representative dataset for this purpose is MT-Bench [62], which evaluates instruction-following ability across eight categories: writing, roleplay, extraction, reasoning, math, coding, STEM, and humanities. The second method involves comparing the model’s generated responses to Davinci-003 responses, using a well-known dataset called Alpaca-Eval [26]. This dataset uses “win_rate” as the evaluation metric.
@@
We choose the LLaMa-7b model as the foundational model for our experiments. To ensure a fair comparison, we also set the maximum input length for the models fine-tuned with the Alpaca dataset to be consistent with LLaMa, which is 2048. In the model fine-tuning phase, we employ the Adam optimizer with a learning rate of $2\times 10^{-5}$ and use a batch size of 64, and conduct training over three epochs. As for the model evaluation phase, we retained all the parameter settings from the original work [26, 62].
@@ 4.2 Experimental Results
The Alpaca dataset encompasses a total of 52,002 instruction examples, with the distribution of their golden scores illustrated in Figure 3. Out of these, 42,460 examples have a golden score exceeding 0.5. We have also chosen some examples that are closely aligned with the golden instructions, notably those with golden scores above 0.8 and 0.85. There are 7,525 examples with golden scores surpassing 0.8 and 619 with golden scores exceeding 0.85. The latter constitutes a mere 1% of the entire dataset.
@@
We conduct instruction tuning on LLaMA using different subsets of examples with golden scores: less than 0.5, greater than 0.5, greater than 0.8, greater than 0.85, and the full dataset. The fine-tuned models are named $\text{Alpaca}_{\leq 0.5}$, $\text{Alpaca}_{>0.5}$, $\text{Alpaca}_{>0.8}$, $\text{Alpaca}_{>0.85}$, and $\text{Alpaca}_{\text{full}}$ respectively.
@@
Table 1: Experimental results on Alpaca-Eval Benchmark.
Model Nums Helpful_Base Koala Self-instruct Oasst Vicuna Length Overall
LLaMA - 0.00 1.28 1.19 0.53 1.25 2,980 0.87
@@
Alpaca${}_{>0.5}$ 42,460 24.03 20.51 28.57 29.78 15.00 413 26.06
Alpaca${}_{>0.8}$ 7,525 34.10 30.76 30.95 35.10 30.00 519 32.48
Alpaca${}_{>0.85}$ 619 37.20 26.90 25.00 29.30 22.50 617 28.20
Table 2: Experimental results on MT-Bench Benchmark.
LLaMA 4.6 4.5 5.2 1.0 1.20 2.2 5.0 4.1 3.47
[…]
Alpaca${}_{>0.5}$ 8.3 5.7 3.5 1.1 1.7 5.0 6.6 7 4.86
Alpaca${}_{>0.8}$ 8.3 5.9 5.6 1.8 2.5 4.0 7.3 7.4 5.34
Alpaca${}_{>0.85}$ 6.6 6.3 4.9 1.0 2.3 3.3 6.3 7.3 4.87
Main Results
The experimental results are presented in Table 1 and Table 2. As expected, $\text{Alpaca}_{>0.8}$ produces the most impressive outcomes. This can be attributed to its ability to maintain an optimal balance between the volume and quality of the instructions it utilizes, leading to the most desirable results. We also note that incorporating lower-quality instructions adversely affected model fine-tuning. This trend is clear when we see that $\text{Alpaca}_{\leq 0.5}$ lagged behind $\text{Alpaca}_{\text{full}}$ in performance, while $\text{Alpaca}_{>0.5}$ shows a slight edge over $\text{Alpaca}_{\text{full}}$. Remarkably, $\text{Alpaca}_{>0.85}$, using only 1% of the dataset for fine-tuning, achieved results comparable to or even surpassing those of $\text{Alpaca}_{\text{full}}$. This underscores the efficacy of our data selection method.
Ablation on Predefined Task Sets
To evaluate how different predefined task sets affect the selection of instruction data for fine-tuning, we include two additional predefined task set variations. One is randomly exampled from the Alpaca dataset but with a smaller task set size, which is limited to 100 examples. The other one entails clustering the Alpaca dataset into 100 clusters using the KMeans algorithm and selecting the centroids of each cluster as examples of the task set.
We use the two predefined sets to calculate golden scores for the Alpaca dataset separately. The distribution of golden scores is depicted in Figure 4. We select instruction data with golden scores less than or equal to 0.5, greater than 0.5, greater than 0.6, greater than 0.7, greater than 0.8, greater than 0.85, and greater than 0.9 for model fine-tuning, respectively. The results in Table 3 suggest that with random sampling, increasing the size of the task set can enhance the identification of high-quality instruction data. The logic behind this is that a larger encompasses a broader diversity of data, facilitating a more nuanced assessment of an instruction’s effect on model task generalization. However, a shift occurs when KMeans is employed to cherry-pick more distinct examples for the task set. With as few as 100 examples, KMeans outshines the results from 1,000 examples acquired through random sampling. In this instance, $\text{Alpaca}_{>0.8}$ delivered a superior performance with just 5,419 examples, compared to the 7,524 examples seen with $\text{Random}_{1000}$. This outcome also indirectly confirms the validity of our hypothesis regarding the definition of golden instructions.
Table 3: Experimental results on Alpaca-Eval Benchmark across different predefined task sets
[…]
Ablation on Instruction Set
To delve deeper into the generalization capabilities of Nuggets across varied directive datasets, we undertake a series of experiments utilizing the Alpaca-GPT4 dataset [34]. It generates instructional data from the powerful GPT-4 model [32], which is considered to have superior data quality. Additionally, it shares the same questions in instructions with the Alpaca dataset, which facilitates our direct comparison between the two.
Inspired by Table 3, we employ the KMeans algorithm on the Alpaca-GPT4 dataset to sample 100 examples, forming the predefined task set. Subsequently, we apply the Nuggets method to score all instructions in the Alpaca-GPT4 dataset with the golden score, as depicted in Figure 5. Compared to the Alpaca dataset, the Alpaca-GPT4 dataset boasts a significantly higher number of instructions with golden scores: 25,100 instructions exceed a score of 0.8, 16,943 surpass 0.85, and 4,250 instructions exceed 0.9. These numbers far exceed the corresponding high-scoring instructions in the Alpaca dataset. This also demonstrates that the golden score can serve as an absolute metric to assess the quality of instructional data. The results from model fine-tuning indicate that on the Alpaca-GPT4 dataset, conclusions align with those of previous experiments. The large language models fine-tuned on subsets with golden scores less than or equal to 0.5 exhibit the poorest performance, with a win rate of only 19.23% in the Alpaca-Eval benchmark. In contrast, the models fine-tuned on subsets with golden scores greater than 0.85 demonstrate superior performance, boasting a high win rate of 72.05%. This success can be attributed to the dual assurance of quantity and quality in this particular subset. It is worth emphasizing that fine-tuning on a small and high-quality dataset consistently and significantly outperforms the results of fine-tuning on the full dataset. Overall, the models fine-tuned using Alpaca-GPT4 significantly outperform those fine-tuned with Alpaca. This implicitly corroborates the superior quality of the Alpaca-GPT4 dataset compared to the Alpaca dataset.
Figure 5: The distribution of the golden score for the instruction dataset across different instruction sets, along with the corresponding fine-tuning results on the Alpaca-Eval benchmark. Both predefined task sets utilize KMeans to sample 100 examples from their respective instruction datasets.
Case Study
To evaluate Nuggets qualitatively, we also select some instruction examples for case studies, as shown in Figure 6. We observe that noisy instructions without outputs display extremely low golden score values. Such examples consistently score very low throughout the instruction dataset. Additionally, instructions with hallucinatory outputs and some that are a bit confusing also typically exhibit lower golden scores. In contrast, instructions with higher scores often possess fluent language and meticulous logic, and are oriented towards assisting humans in problem-solving. Therefore, we believe that the high golden scores can represent high-quality instruction tuning data.
5 Discussion: One Shot Learning as Implicit Instruction Tuning
A widely acknowledged philosophy is that instruction tuning significantly enhances the task generalization of pre-trained models to perform various specific tasks [31, 57, 58, 38]. Therefore, we introduce the hypothesis of a golden instruction: the evaluation of an instruction’s effectiveness is determined by its impact on the model’s task generalization capability following the fine-tuning process. As the degree of improvement becomes more pronounced, the instruction approaches a classification as “golden instruction”.
Based on this hypothesis, a straightforward method is to fine-tune an independent model using an instruction example and then compare the performance of the fine-tuned model with the base model on a predefined dataset containing multiple tasks, to determine whether this example qualifies as a golden instruction. However, this way will result in an impractical number of fine-tuned models, equivalent to the number of instructions. Additionally, fine-tuning with only one example may introduce unstable updates to the model’s gradients, making it challenging to determine whether the example has genuinely been learned. Motivated by the inherent duality between attention and gradient descent [11, 2, 55, 18], we incorporate the meta-optimize mechanism into Transformer attention, replacing the need for actually fine-tuning the model.
[…]
@@ 5 Discussion: One Shot Learning as Implicit Instruction Tuning
where $\sqrt{d_{\text{in}}}$ serves as a scaling factor. The term $\mathbf{W}_{V}\mathbf{X}_{\text{test}}(\mathbf{W}_{K}\mathbf{X}_{\text{test}})^{\top}$ could be denoted as $\mathbf{W}_{\text{zsl}}$, which represents the zero shot learning scenario where no instruction tuning is performed, since it solely focuses on the test input. In addition, the term $\mathbf{W}_{V}\mathbf{X}_{\text{ins}}(\mathbf{W}_{K}\mathbf{X}_{\text{ins}})^{\top}$ can be seen as implicit instruction tuning $\Delta\mathbf{W}_{\text{iit}}$ achieved via the meta-gradient [11, 55] derived from the instruction sample. Readers can refer to previous papers [11, 2, 18] for more details on implicit instruction tuning.
@@ 6 Conclusion
In this paper, we introduce Nuggets, a method that utilizes large language models to uncover more critical data for instruction tuning. It is based on one shot learning to detect the value of examples, which allows for efficient data selection and does not rely on additional annotation and cost. Benefiting from Nuggets, we find that even smaller subsets yield better instruction followings. Furthermore, we argue that our method emphasizes the need for data selection and provides insights for future instruction fine-tuning processes, encouraging fine-grained data combinations rather than blindly expanding quantities.
Future Work
While its efficacy has been confirmed in our experiments, there remain avenues for improvement. First, we can investigate including a diverse and compact task set in the golden scoring phase. The investigation will facilitate a more efficient model evaluation of instructional data and better identification of high-quality instructions for model fine-tuning. Second, due to resource constraints, the experiments of this work are limited to the LLaMA-7b model. While it is one of the most influential works in the large language model open-source community, further validation across a wider range of models is essential. Lastly, it is important to validate the proposed method on a more extensive set of instructional datasets.

### v4

@@ 4.1 Experimental Setup
We adopt the Alpaca dataset Taori et al. (2023) as instruction data. It is an important resource in the open-source community for instruction tuning, which is constructed by employing the self-instruct Wang et al. (2022a) method to distill instruction data from text-davinci-003. The success of this dataset in fine-tuning the LLaMA model has sparked a series of explorations into instruction fine-tuning Li et al. (2023b); Ji et al. (2023); Xu et al. (2023b). Besides, we perform more types of instruction datasets to verify the transferability of Nuggets, please refer to B.
@@
The predefined task set plays a crucial role in computing golden scores for instructions. These data are employed to evaluate the model’s ability to generalize across diverse tasks. The adequacy of the predefined task set is contingent upon its encompassing a substantial volume of data and incorporating a broad range of tasks. As the Alpaca dataset inherently possesses these attributes, we randomly choose 1,000 examples from it to constitute the predefined task set.
@@
This work uses two methods to assess the model’s capabilities. The first approach involves rating the responses generated by models on a scale ranging from 1 to 10. For this purpose, we utilize the GPT-4 labeled MT-Bench Zheng et al. (2023) dataset, which evaluates instruction-following proficiency across eight categories: writing, roleplay, extraction, reasoning, math, coding, STEM, and humanities. Notably, since we only fine-tune on single-turn instruction data, the evaluation is restricted to Turn 1 of MT-Bench, similar to previous studies Cao et al. (2023); Zheng et al. (2023); Chen et al. (2023). The second method involves comparing the model’s generated responses with those produced by the Davinci-003 model, employing the well-established Alpaca-Eval dataset Li et al. (2023d). This dataset adopts the “win_rate” as the evaluation metric.
@@
In our experiments, we designate the LLaMA-7B model as the foundational model. To ensure a fair comparison, we also set the maximum input length for the models fine-tuned with the Alpaca dataset to be consistent with LLaMA, which is 2048. In the model fine-tuning phase, we employ the Adam optimizer with a learning rate of $2\times 10^{-5}$ and utilize a batch size of 64, conducting training over three epochs. In the subsequent model evaluation phase, we maintain all parameter settings consistent with the original work Li et al. (2023d); Zheng et al. (2023).
@@ 4.2 Experimental Results
The Alpaca dataset comprises a total of 52,002 instruction examples, and the distribution of their golden scores is illustrated in Appendix A. Among these examples, 42,460 instances exhibit a golden score surpassing 0.5. In addition, a subset of examples closely aligned with the golden instructions has been selected, specifically those attaining golden scores above 0.8 and 0.85. In particular, there are 7,525 examples with golden scores surpassing 0.8 and 619 examples with golden scores exceeding 0.85. Notably, the latter subset constitutes a mere 1% of the entire dataset.
@@
We conduct instruction tuning on the LLaMA model using various subsets of examples distinguished by their golden scores: those with scores less than 0.5, greater than 0.5, greater than 0.8, greater than 0.85, and the complete dataset. The fine-tuned models are denoted as $\text{Alpaca}_{\leq 0.5}$, $\text{Alpaca}_{>0.5}$, $\text{Alpaca}_{>0.8}$, $\text{Alpaca}_{>0.85}$, and $\text{Alpaca}_{\text{full}}$, respectively.
Predefined Task Set Alpaca≤0.5 Alpaca>0.5 Alpaca>0.6 Alpaca>0.7 Alpaca>0.8 Alpaca>0.85 Alpaca>0.9
K-Means100 11.91 24.44 23.94 25.93 34.25 25.25 17.35
Random100 9.65 22.28 24.16 26.56 31.67 27.74 26.34
Random1000 10.96 26.06 24.46 28.43 32.48 28.21 -
@@
Table 3: Win_rate results on Alpaca-Eval Benchmark across different predefined task sets
Main Results
The experimental results are presented in Table 1 and Table 2 for the Alpaca-Eval and MT-Bench benchmarks, respectively. As expected, $\text{Alpaca}_{>0.8}$ produces the most impressive outcomes. This can be attributed to its ability to maintain an optimal balance between the volume and quality of the instructions it utilizes, leading to the most desirable results. We also note that incorporating lower-quality instructions adversely affected model fine-tuning. This trend is clear when we see that $\text{Alpaca}_{\leq 0.5}$ lagged behind $\text{Alpaca}_{\text{full}}$ in performance, while $\text{Alpaca}_{>0.5}$ shows a slight edge over $\text{Alpaca}_{\text{full}}$. Remarkably, $\text{Alpaca}_{>0.85}$, using only 1% of the dataset for fine-tuning, achieved results comparable to or even surpassing those of $\text{Alpaca}_{\text{full}}$. This underscores the efficacy of our data selection method. More qualitative results can be found in Appendix C.
Figure 4: The distribution of the golden score for the instruction dataset across different instruction sets, along with the corresponding fine-tuning results on the Alpaca-Eval benchmark. Both predefined task sets utilize K-Means to sample 100 examples from their respective instruction datasets.
GS≤0.5 GS>0.5 GS>0.6 GS>0.7 GS>0.8 GS>0.85 GS>0.9 Full Data
LLaMA2 NUM 3,730 48,272 40,905 28,644 10,409 2,411 87 52,002
Win_rate 13.17 27.09 27.85 27.62 33.92 34.98 27.08 26.47
Mistral NUM 78 51,924 51,610 49,398 36,068 23,147 9,356 52,002
Win_rate 0 12.26 11.10 12.45 11.28 10.60 13.53 9.85
Table 4: Win_rate results on Alpaca-Eval Benchmark across two different foundation models.
Ablation on Predefined Task Sets
To evaluate how different predefined task sets affect the selection of instruction data for fine-tuning, we include two additional predefined task set variations. One is randomly exampled from the Alpaca dataset but with a smaller task set size, which is limited to 100 examples. The other one entails clustering the Alpaca dataset into 100 clusters using the K-Means algorithm and selecting the centroids of each cluster as examples of the task set.
We use the two predefined sets to calculate golden scores for the Alpaca dataset separately. The distribution of golden scores is depicted in Figure 3. We select instruction data with golden scores less than or equal to 0.5, greater than 0.5, greater than 0.6, greater than 0.7, greater than 0.8, greater than 0.85, and greater than 0.9 for model fine-tuning, respectively. Table 3 suggests that with random sampling, increasing the size of the task set can enhance the identification of high-quality instruction data. The logic behind this is that a larger encompasses a broader diversity of data, facilitating a more nuanced assessment of an instruction’s effect on model task generalization. However, a shift occurs when K-Means is employed to cherry-pick more distinct examples for the task set. With as few as 100 examples, K-Means outshines the results from 1,000 examples acquired through random sampling. In this instance, $\text{Alpaca}_{>0.8}$ delivered a superior performance with just 5,419 examples, compared to the 7,524 examples seen with $\text{Random}_{1000}$. This outcome also indirectly confirms the validity of our hypothesis regarding the definition of golden instructions.
Ablation on Instruction Sets
To delve deeper into the generalization capabilities of Nuggets across varied instruction datasets, we undertake a series of experiments utilizing the Alpaca-GPT4 dataset Peng et al. (2023). It generates instructional data from the powerful GPT-4 model OpenAI (2023), which is considered to have superior data quality. Additionally, it shares the same questions in instructions with the Alpaca dataset, which facilitates our direct comparison between the two.
Inspired by Table 3, we employ the K-Means algorithm on the Alpaca-GPT4 dataset to sample 100 examples, forming the predefined task set. Subsequently, we apply the Nuggets method to score all instructions in the dataset with the golden score, as depicted in Figure 4. Compared to the Alpaca dataset, the Alpaca-GPT4 dataset boasts a higher number of instructions with golden scores: 25,100 instructions exceed a score of 0.8, 16,943 surpass 0.85, and 4,250 instructions exceed 0.9. These numbers far exceed the corresponding high-scoring instructions in the Alpaca dataset. This also demonstrates that the golden score can serve as an absolute metric to assess the quality of instructional data. The results from model fine-tuning indicate that on the Alpaca-GPT4 dataset, conclusions align with those of previous experiments. The large language models fine-tuned on subsets with golden scores less than or equal to 0.5 exhibit the poorest performance, with a win rate of only 19.23% in the Alpaca-Eval benchmark. In contrast, the models fine-tuned on subsets with golden scores greater than 0.85 demonstrate superior performance, boasting a high win rate of 72.05%. This success can be attributed to the dual assurance of quantity and quality in this particular subset. It is worth emphasizing that fine-tuning on a small and high-quality dataset consistently and significantly outperforms the results of fine-tuning on the full dataset. Overall, the models fine-tuned using Alpaca-GPT4 significantly outperform those fine-tuned with Alpaca. This implicitly corroborates the superior quality of the Alpaca-GPT4 dataset compared to the Alpaca dataset. For more experiments on instruction datasets, please refer to Appendix B.
Ablation on Foundation Models
To verify the transferability of the Nuggets method, we conducted experiments on different foundation models using the Alpaca instruction dataset. We selected LLaMA2 (Touvron et al., 2023b) and Mistral (Jiang et al., 2023) at the 7B size as the new base models. The distribution of the golden scores and the performance of models fine-tuned on corresponding subsets of instructions are shown in Table 4. We found that the Nuggets method is also applicable to other models. LLaMA2 achieved the best results under fine-tuning on subsets with a golden score greater than 0.85, reaching 34.98, which is significantly higher than the 26.47 achieved under full data. Although the absolute value of the win_rate for the Mistral series of fine-tuned models is somewhat low, their performance is also significantly boosted by the Nuggets data filtering.
5 Discussion: One-Shot Learning as Implicit Instruction Tuning
@@
$=\mathbf{W}_{V}[\mathbf{X}_{\text{ins}}\|\mathbf{X}_{\text{test}}]\textsf{Softmax}\left(\frac{\mathbf{W}_{K}[\mathbf{X}_{\text{ins}}\|\mathbf{X}_{\text{test}}]^{\top}\textit{{Q}}}{\sqrt{d_{\text{in}}}}\right)$
@@
where $\sqrt{d_{\text{in}}}$ serves as a scaling factor. The term $\mathbf{W}_{V}\mathbf{X}_{\text{test}}(\mathbf{W}_{K}\mathbf{X}_{\text{test}})^{\top}$ could be denoted as $\mathbf{W}_{\text{zsl}}$, which represents the zero-shot learning scenario where no instruction tuning is performed since it solely focuses on the test input. In addition, the term $\mathbf{W}_{V}\mathbf{X}_{\text{ins}}(\mathbf{W}_{K}\mathbf{X}_{\text{ins}})^{\top}$ can be seen as implicit instruction tuning $\Delta\mathbf{W}_{\text{iit}}$ achieved via the meta-gradient Dai et al. (2022); Yang et al. (2023) derived from the instruction sample. Readers can refer to previous papers Dai et al. (2022); Aizerman et al. (1964); Irie et al. (2022) for more details on implicit instruction tuning.
@@ 6 Conclusion
This paper presents Nuggets, a method leveraging LLMs to discern more pivotal data for instruction tuning. Grounded in one-shot learning, this approach facilitates the identification of examples’ value, enabling efficient data selection without dependence on additional annotation and associated costs. Benefiting from Nuggets, we observe improved instruction following abilities even with smaller training subsets. Furthermore, we posit that our method underscores the significance of meticulous data selection, offering valuable insights for future instruction fine-tuning endeavors.
Limitations
Although the efficacy of the proposed approach has been confirmed through empirical experiments, opportunities for refinement persist. One avenue for improvement involves a thorough investigation into the inclusion of a diverse and compact set of predefined tasks during the golden scoring phase. This exploration aims to enhance the efficiency of model evaluation on instructional data, leading to improved identification of high-quality instructions suitable for subsequent model fine-tuning. Secondly, due to resource constraints, the majority of experiments in this study are confined to the LLaMA-7B model. While this model holds significant influence within the large language model open-source community, comprehensive validation across a broader spectrum of models is imperative to ensure the generalizability of the proposed approach. Lastly, to fortify the empirical foundation of our findings, it is crucial to subject the proposed method to validation on a more extensive array of instructional datasets. This step aims to ascertain the robustness and applicability of the methodology across a diverse range of instructional contexts, contributing to its broader utility in real-world scenarios. These outlined avenues for future work are anticipated to refine and extend the scope of our proposed method.
Acknowledgments
Min Yang was supported by National Key Research and Development Program of China (2022YFF0902100), National Natural Science Foundation of China (62376262), the Natural Science Foundation of Guangdong Province of China (2024A1515030166), Shenzhen Science and Technology Innovation Program (KQTD20190929172835662), Shenzhen Basic Research Foundation (JCYJ20210324115614039). This work was supported by Alibaba Group through Alibaba Innovative Research Program. Xiaobo Xia was supported by the Australian Research Council project: DE190101473 and Google PhD Fellowship. Tongliang Liu is partially supported by the following Australian Research Council projects: FT220100318, DP220102121, LP220100527, LP220200949, and IC190100031.
Appendix A The Distribution of Golden Score
As shown in Figure 5, in a total of 52,002 cases, there are 9,549 instructions with a gold score of less than 0.5, indicating that these data have a side effect on overall task completion. Besides, there are 7,524 instructions with a gold score greater than 0.8, suggesting that the model improves the task completion rate through one-shot learning from these data, which can be considered high-quality instruction data.
Figure 5: The distribution of the golden score for the Alpaca instruction dataset.
Appendix B Experiment on Other Instruction Sets
Based on the LLaMA-7B model, we conducted experiments on several other instruction datasets, further validating the effectiveness of our Nuggets method.
B.1 Code Alpaca
The Code Alpaca instruction dataset (Chaudhary, 2023) is designed to develop large language models capable of following instructions and generating code. Leveraging self-instruct (Wang et al., 2022a) technology, it has produced 20,000 examples of instruction data. We use HumanEval (Chen et al., 2021) as a benchmark to evaluate the model’s code generation capabilities. It is used to measure functional correctness for synthesizing programs from docstrings. It consists of 164 original programming problems, assessing language comprehension, algorithms, and simple mathematics, with some being comparable to simple software interview questions. We adopt the approach outlined by Chen et al. (2021) to calculate pass rates at k values of 1, 10, and 100 for each problem. Essentially, pass@1 predicts the probability of a model producing a correct solution on the first try, while pass@10 and pass@100 predict the probability of achieving a correct solution within 10 and 100 tries, respectively. We generate 200 completions at a temperature setting of 0.2 (Luo et al., 2023) to estimate pass@1, pass@10, and pass@100 rates.
Figure 6: The distribution of the golden score for the Code Alpaca instruction dataset, along with the corresponding fine-tuning results on the HumanEval benchmark. Predefined task sets utilize K-Means to sample 100 examples from the Code Alpaca instruction dataset.
The experimental results are shown in the Figure 6. Out of 20,000 instructions, 4,715 instructions have a gold score greater than 0.85, achieving the best pass@1 and pass@10 results in the HumanEval benchmark, superior to the fine-tuning results of the full dataset. Additionally, this experiment also proves that the Nuggets method can be applied to fine-tuning for specific tasks, demonstrating good transferability.
B.2 WizardLM
The WizardLM instruction dataset Xu et al. (2023a), which employs Evol-Instruct to iteratively refine an initial set of instructions into more complex ones, contains 70,000 instruction examples. The distribution of the golden scores and the performance of models fine-tuned on corresponding subsets of instructions are shown in Table 5.
GS≤0.5 GS>0.5 GS>0.6 GS>0.7 GS>0.8 GS>0.85 GS>0.86 GS>0.87 Full Data
NUM 480 69,520 69,377 68,898 65,190 40,223 23,579 3, 316 70,000
Win_rate 19.42 58.08 57.40 56.21 59.81 58.40 57.81 54.68 57.65
Table 5: The distribution of golden scores for the WizardLM dataset and the evaluation results of models fine-tuned on corresponding score subsets on the Alpaca-Eval benchmark.
We can observe that the quality distribution of the WizardLM dataset is relatively balanced, with 65,190 instruction examples having a golden score greater than 0.8, accounting for 93% of the total number of instructions. In the evaluation of the Alpaca-Eval benchmark, models fine-tuned on subsets with golden scores greater than 0.8 achieved a win rate of 59.81, outperforming models fine-tuned on the full dataset.
B.3 FLANv2
We sampled 50,000 examples from the FLANv2 (Chung et al., 2022) dataset to constitute the instruction tuning data for this experiment. Additionally, the Predefined task set was also derived from these 50,000 examples, using the K-Means algorithm to sample 100 examples. We evaluated the performance of the fine-tuned model using MMLU Hendrycks et al. (2020) in a 5-shot setting. MMLU is a test designed to measure a text model’s multitask accuracy. The test encompasses 57 tasks, including elementary mathematics, US history, computer science, law, and more.
GS≤0.5 GS>0.5 GS>0.6 GS>0.7 GS>0.8 GS>0.85 GS>0.9 Full Data
NUM 1,361 48,639 46,009 39,046 17,037 4,798 321 50,000
Acc 34.68 41.38 41.92 41.87 41.97 35.51 26.41 40.45
Table 6: The distribution of golden scores for the sampled FLANv2 dataset and the evaluation results of models fine-tuned on corresponding score subsets on the MMLU benchmark.
The experimental results are shown in the table. It can be observed that the model fine-tuned with examples having a golden score greater than 0.8 (totaling 17,037 examples) achieved the best results, followed by those fine-tuned with examples having a golden score greater than 0.6 (totaling 46,009 examples). Additionally, we noted that the model fine-tuned with examples having a golden score less than 0.5 (totaling 1,361 examples) outperformed the model fine-tuned with examples having a golden score greater than 0.9 (totaling 321 examples).
We speculate that the reason for this outcome is that the MMLU evaluation itself is a knowledge-intensive assessment, which thus imposes higher demands on the quantity of fine-tuning examples. In the future, some optimizations in the selection of the Predefined Task Set may potentially improve this situation.
Appendix C Case Study
To evaluate Nuggets qualitatively, we also select some instruction examples from the Alpaca dataset for case studies, as shown in Figure 7. We observe that noisy instructions without outputs display extremely low golden score values. Such examples consistently score very low throughout the instruction dataset. Additionally, instructions with hallucinatory outputs and some that are a bit confusing also typically exhibit lower golden scores. In contrast, instructions with higher scores often possess fluent language and meticulous logic, and are oriented towards assisting humans in problem-solving. Therefore, we believe that the high golden scores can represent high-quality instruction tuning data.